If a circular disc is measured as 2.12cm using a device of least count 0.1 mm.Find the possible error of the area of the disc?
1
Expert's answer
2020-05-08T16:12:35-0400
"\frac{\Delta A}{A}=2\frac{\Delta r}{r}=2\frac{0.1}{21.2}"

"\Delta A=2\frac{0.1}{21.2}\pi(21.2)^2=10\ mm^2"